# Attribute-driven Disentangled Representation Learning for Multimodal Recommendation

### v1

@@ Abstract.
Recommendation algorithms forecast user preferences by correlating user and item representations derived from historical interaction patterns. In pursuit of enhanced performance, many methods focus on learning robust and independent representations by disentangling the intricate factors within interaction data across various modalities in an unsupervised manner. However, such an approach obfuscates the discernment of how specific factors (e.g., category or brand) influence the outcomes, making it challenging to regulate their effects. In response to this challenge, we introduce a novel method called Attribute-Driven Disentangled Representation Learning (short for AD-DRL), which explicitly incorporates attributes from different modalities into the disentangled representation learning process. By assigning a specific attribute to each factor in multimodal features, AD-DRL can disentangle the factors at both attribute and attribute-value levels. To obtain robust and independent representations for each factor associated with a specific attribute, we first disentangle the representations of features both within and across different modalities. Moreover, we further enhance the robustness of the representations by fusing the multimodal features of the same factor. Empirical evaluations conducted on three public real-world datasets substantiate the effectiveness of AD-DRL, as well as its interpretability and controllability.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recommender systems (RS) are integral to a myriad of online platforms, spanning E-commerce to advertising, facilitating users in pinpointing items aligned with their preferences. Given their pivotal role, numerous efforts have been dedicated to developing advanced recommendation models to improve their performance. Among them, Collaborative Filtering (CF) based models (Rendle et al., 2009; Rendle, 2010; He et al., 2017; Hsieh et al., 2017; Wang et al., 2019; He et al., 2020; Wang et al., 2020) have achieved great success by exploiting the user-item interaction data to learn user and item representation. However, these models can easily encounter the sparsity problem in practical scenarios due to their exclusive dependence on interaction data. To alleviate the data sparsity problem in recommendation, side information (e.g., attributes, user reviews and item images), which contains rich information associated with users or items, is often used to enhance the representation learning of users and items (Zhang et al., 2016, 2017a; Liu et al., 2019; Wei et al., 2019, 2020; Liu et al., 2022, 2023).
@@
It is well recognized that the entangled representations of users and items are infeasible to directly capture fine-grained user preferences across diverse factors, thereby constraining both the efficacy and interpretability of recommender systems (Ma et al., 2019; Wang et al., 2020). In recent years, the study of disentangled representation learning has garnered significant attention in diverse fields, notably in computer vision (Higgins et al., 2017; Liu et al., 2020; Hsieh et al., 2018), due to its capability to identify and disentangle the underlying factors behind data. Empirical evidence suggests that disentangled representations exhibit enhanced robustness, particularly in intricate application contexts. Hence, many recommendation methods adopt disentangled representation learning techniques to learn robust and independent representations, ultimately enhancing recommendation performance (Ma et al., 2019; Mu et al., 2022; Wang et al., 2020; Tran and Lauw, 2022; Liu et al., 2022). For example, Ma et al. (Ma et al., 2019) employed disentangled representations to capture user preferences regarding different concepts associated with user intentions. Wang et al. (Wang et al., 2020) introduced a GCN-based model that produces disentangled representations by modelling a distribution over intents for each user-item interaction, exploring the diversity of user intentions on adopting items. However, these methods only concentrate on disentangling the user and item representations based on their ID embeddings. To exploit the difference between the factors behind data of various modalities, Liu et al. (Liu et al., 2022) estimated the users’ attention weight to underlying factors of different modalities, utilizing a sophisticated attention-driven module.
@@
Despite the considerable advancements brought about by disentangling techniques in recommendation systems, existing studies often disentangle both users and items into latent factors. This methodology inherently constrains the model’s interpretability and controllability. For example, consider using disentangled representations to elucidate the diverse factors, such as style, brand, popularity, and price, that shape user preferences in dress selection. The inherent abstraction of latent factors from current disentangling methods makes it arduous to pinpoint which factor represents each specific dress attribute. In particular, one factor might be loosely related to a combination of brand and price, while another factor might represent a mixture of style and popularity. This ambiguity in the latent factors limits the interpretability of the recommendation system, making it difficult to understand why a particular item was recommended to a user. Moreover, this lack of clarity also affects the controllability of the recommendation system. Suppose a user wants to receive recommendations specifically for dresses without considering her preferences for price and popularity. It would be challenging to adjust the recommendation system to focus on those specific preferences, as the latent factors are not clearly tied to these attributes.
@@
In summary, the contribution of this paper is threefold:
@@ 2. Related Work
In this section, we present a concise overview of the latest advancements in multimodal collaborative filtering techniques and disentangled representation learning in recommendation, as they hold considerable pertinence to our study.
@@ 2.1. Multimodal Collaborative Filtering
Traditional Collaborative Filtering (CF) (Sarwar et al., 2001; Ebesu et al., 2018; Rendle et al., 2009) methods primarily rely on user-item interaction behaviour to learn representations of users and items. Consequently, the recommendation performance is negatively impacted when encountering users and items with limited interactions. To mitigate the challenges posed by data sparsity (McAuley and Leskovec, 2013; Kalantidis et al., 2013; Tan et al., 2016), recent research has incorporated multimodal information into recommendation systems. The multimodal information, such as reviews and images, provides valuable information of user preference and item characteristics, which can supplement historical user-item interactions and thus enhance recommendation performance (He and McAuley, 2016; Zhang et al., 2017a; Liu et al., 2019; Wei et al., 2019, 2020; Liu et al., 2022, 2023).
Early CF models like matrix factorization (MF) (Koren et al., 2009) project the IDs of users and items to embeddings by minimizing the error of reconstructing the user-item interaction matrix. Previous studies integrate multimodal information into the MF-based method in a straightforward manner. For instance, VBPR (He and McAuley, 2016) directly concatenates the visual features learned from item images with collaborative features as the joint item representation and feeds it into the MF module. With the success of deep learning techniques in modeling complex interaction behaviors (He et al., 2017) and the relation between various multimodal features (Nie et al., 2016), a lot of deep learning techniques have been introduced into multimodal recommendation (He et al., 2017; Tay et al., 2018). For example, MAML (Liu et al., 2019) first fuses the item’s multimodal features and user features, then feeds it into an attention module to capture users’ diverse preferences. VECF (Chen et al., 2019) constructs a visually explainable collaborative filtering model using a multimodal attention network, facilitating the integrated coupling of diverse feature modalities. More recently, graph convolutional networks (GCNs) (Hamilton et al., 2017) have demonstrated their power capability of representation learning for recommendation (Wang et al., 2019; He et al., 2020; Zhou et al., 2023; Wei et al., 2023). Based on the GCN structure, MMGCN (Wei et al., 2019) constructs a user-item bipartite graph to learn representations of each modality and then fuse the modality representation together as the final representation. As there are false-positive edges exist which will hurt the performance of recommendation, GRCN (Wei et al., 2020) exploits the rich multimodal content of items to refine the structure of the interaction graph.
@@ 2.2. Disentangled Representation Learning
Disentangled representation learning, which seeks to identify and separate underlying explanatory factors embedded in data, has garnered significant interest, particularly within the realm of computer vision (Higgins et al., 2017; Liu et al., 2020; Ge et al., 2021). For instance, the beta-VAE method (Higgins et al., 2017) employs a constrained variational framework to learn disentangled representations of fundamental visual concepts. Meanwhile, IPGDN (Liu et al., 2020) automatically uncovers independent latent factors present in graph data.
Owing to the successful application of disentangled representation learning in various domains, numerous studies have concentrated on learning disentangled representations for users and items in recommendation systems in recent years (Ma et al., 2019; Tran and Lauw, 2022; Wang et al., 2020; Mu et al., 2022; Liu et al., 2022; Wang et al., 2023). Following the previous work in computer vision, the initial attempts take Variational Auto-encoder (VAE) (Kingma and Welling, 2014) to learn disentangled representations. For example, MacridVAE (Ma et al., 2019) captured user preferences regarding the different concepts associated with user intentions separately. Beyond the disentangled representations derived from user-item interaction modeling, ADDVAE (Tran and Lauw, 2022) acquires an additional set of disentangled user representations from textual content and subsequently aligns both sets of representations. For studying the diversity of user intents on adopting the items, DGCF (Wang et al., 2020) employs a graph disentangled module to iteratively refine the intent-aware interaction graph and factorial representations for recommendations. In order to model the users’ various preferences on different factors of each modality, DRML (Liu et al., 2022) estimates the user’s attention weight to underlying factors of different modalities with an attention module. As the learned disentangled representations lack clear meanings, KDR (Mu et al., 2022) harnesses the power of knowledge graphs (KGs) to guide the disentangled representation learning process, ensuring that the resulting disentangled representations are associated with semantically meaningful information extracted from the KGs.
@@ 3. Method
In this section, we first formally formulate the problem statement of multimodal recommendation and disentangled representation learning in Section 3.1. We then detail our designed attribute-driven disentangled representation learning in Section 3.2. Lastly, we describe the preference prediction process of our model in Section 3.3.
@@ 3.1. Preliminaries
Before delving into the details of our model, we first outline the problem setting and the intuition of attribute-driven disentanglement.
@@ 3.1.1. Problem Setting
Given a set of users $u\in\{1,2,\cdots,N_{u}\}$ and items $i\in\{1,2,\cdots,N_{i}\}$, we utilize two types of information to learn user and item representations: (1) A user-item interaction matrix $\mathbf{R}\in\mathbb{R}^{N_{u}\times N_{i}}$, where each entry $r_{ui}\in\mathbf{R}$ represents the implicit feedback (e.g., clicks, likes or purchases) of user $u$ to item $i$. $r_{ui}=1$ indicates that user $u$ has interacted with item $i$ and $r_{ui}=0$ indicates that there is no interaction between user $u$ and item $i$ in the observed data; and (2) multimodal features, which mainly consist of two types of features associated with items, namely reviews and images. By leveraging both user-item interactions and multimodal features, we aim to learn robust representations of users and items, thereby predicting a user’s preference to an uninteracted item.
@@
Similar to previous work (Wei et al., 2020; Liu et al., 2022), each user $u$ and item $i$ are assigned with a unique ID and respectively represented by a vector $\bm{p_{u}}\in\mathbb{R}^{d}$ and $\bm{q_{i}}\in\mathbb{R}^{d}$, which are randomly initialized in our model. For the review and image information, we use the BERT (Devlin et al., 2019) and the ViT (Dosovitskiy et al., 2021) to extract the raw textual features $\bm{F_{t}}\in\mathbb{R}^{d_{0}}$ and visual features $\bm{F_{v}}\in\mathbb{R}^{d_{0}}$, respectively. To tailor the recommendation-oriented features, we adopt two non-linear transformations to cast $\bm{F_{t}}$ and $\bm{F_{v}}$ into the same feature space:
(1) $\bm{q_{t}}$ $=\sigma(\textbf{W}_{t}\bm{F_{t}}+b_{t}),$
$\bm{q_{v}}$ $=\sigma(\textbf{W}_{v}\bm{F_{v}}+b_{v}),$
where $\textbf{W}_{t}$, $\textbf{W}_{v}\in\mathbb{R}^{d\times d_{0}}$ and $b_{t}$, $b_{v}\in\mathbb{R}^{d}$ denote the weight matrices and bias vectors for textual and visual modality, respectively. $\sigma(\cdot)$ is the activation function.
Following the disentangled representation learning process in DMRL (Liu et al., 2022), we first split the feature vector into $K$ chunks. Each chunk corresponds to a specific item attribute, such as price, brand, etc. For simplicity, we equally split the representation of each modality into $K$ continuous chunks. Take item ID embedding as an example:
(2) $\bm{q_{i}}=(\bm{q_{i}^{1}},\bm{q_{i}^{2}},\cdots,\bm{q_{i}^{K}}),$
where $\bm{q_{i}^{k}}\in\mathbb{R}^{\frac{d}{K}}$ is the item ID corresponding to the $k$-th attribute. Analogously, $\bm{q_{t}}=(\bm{q_{t}^{1}},\bm{q_{t}^{2}},\cdots,\bm{q_{t}^{K}})$, $\bm{q_{v}}=(\bm{q_{v}^{1}},\bm{q_{v}^{2}},\cdots,\bm{q_{v}^{K}})$ and $\bm{p_{u}}=(\bm{p_{u}^{1}},\bm{p_{u}^{2}},\cdots,\bm{p_{u}^{K}})$ are defined as the embeddings of textual feature, visual feature and user ID embedding, respectively.
@@ 3.1.3. Intuition of Attribute-driven Disentanglement
Our work not only aims to recommend accurate items to users but also advocates for attribute-driven disentanglement in multimodal recommender systems to enhance the interpretability and controllability of recommendations. Specifically, unlike the existing method that disentangles the latent factors in an unsupervised manner, we leverage the semantic labels of item attributes to learn attribute-specific sub-spaces for each attribute to obtain disentangled representations. In this paper, the vector of each modality is composed of $K$ chunks, with the assumption that each chunk is associated with an attribute (such as price or brand). The achievement of attribute-driven disentangling requires two necessary conditions. Firstly, each chunk should have a clear attribute reference, and chunks associated with different attributes should be distinguishable. For example, chunk $i$ is associated with the price, while chunk $j$ represents the vector of the brand. Secondly, each chunk of a specific attribute should accurately capture the specific attribute value. For example, chunk $i$ should reflect the price level (i.e., expensive or cheap) of the product.

### v2

@@ Abstract.
Recommendation algorithms predict user preferences by correlating user and item representations derived from historical interaction patterns. In pursuit of enhanced performance, many methods focus on learning robust and independent representations by disentangling the intricate factors within interaction data across various modalities in an unsupervised manner. However, such an approach obfuscates the discernment of how specific factors (e.g., category or brand) influence the outcomes, making it challenging to regulate their effects. In response to this challenge, we introduce a novel method called Attribute-Driven Disentangled Representation Learning (short for AD-DRL), which explicitly incorporates attributes from different modalities into the disentangled representation learning process. By assigning a specific attribute to each factor in multimodal features, AD-DRL can disentangle the factors at both attribute and attribute-value levels. To obtain robust and independent representations for each factor associated with a specific attribute, we first disentangle the representations of features both within and across different modalities. Moreover, we further enhance the robustness of the representations by fusing the multimodal features of the same factor. Empirical evaluations conducted on three public real-world datasets substantiate the effectiveness of AD-DRL, as well as its interpretability and controllability.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681148isbn: 979-8-4007-0686-8/24/10ccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recommender systems (RS) are integral to a myriad of online platforms, spanning E-commerce to advertising, facilitating users in pinpointing items aligned with their preferences. Given their pivotal role, numerous efforts have been dedicated to developing advanced recommendation models to improve their performance. Among them, Collaborative Filtering (CF) based models (Rendle et al., 2009; Rendle, 2010; He et al., 2017; Hsieh et al., 2017; Wang et al., 2019; He et al., 2020; Wang et al., 2020; Liu et al., 2021) have achieved great success by exploiting the user-item interaction data to learn user and item representation. However, these models can easily encounter the sparsity problem in practical scenarios due to their exclusive dependence on interaction data. To alleviate the data sparsity problem in recommendation, side information (e.g., attributes, user reviews and item images), which contains rich information associated with users or items, is often used to enhance the representation learning of users and items (Zhang et al., 2016, 2017a; Liu et al., 2019; Wei et al., 2019, 2020; Liu et al., 2023a, c).
@@
It is well recognized that the entangled representations of users and items are infeasible to directly capture fine-grained user preferences across diverse factors, thereby constraining both the efficacy and interpretability of recommender systems (Ma et al., 2019; Wang et al., 2020). In recent years, the study of disentangled representation learning has garnered significant attention in diverse fields, notably in computer vision (Higgins et al., 2017; Liu et al., 2020; Hsieh et al., 2018), due to its capability to identify and disentangle the underlying factors behind data. Empirical evidence suggests that disentangled representations exhibit enhanced robustness, particularly in complex application contexts. Hence, many recommendation methods adopt disentangled representation learning techniques to learn robust and independent representations, ultimately enhancing recommendation performance (Ma et al., 2019; Mu et al., 2022; Wang et al., 2020; Tran and Lauw, 2022; Liu et al., 2023a). For example, Ma et al. (Ma et al., 2019) employed disentangled representations to capture user preferences regarding different concepts associated with user intentions. Wang et al. (Wang et al., 2020) introduced a GCN-based model that produces disentangled representations by modelling a distribution over intents for each user-item interaction, exploring the diversity of user intentions on adopting items. However, these methods only concentrate on disentangling the user and item representations based on their ID embeddings. To exploit the difference between the factors behind data of various modalities, Liu et al. (Liu et al., 2023a) estimated the users’ attention weight to underlying factors of different modalities, utilizing a sophisticated attention-driven module.
@@
Despite the considerable advancements brought about by disentangling techniques in recommender systems, existing studies often disentangle both users and items into latent factors. This methodology inherently constrains the model’s interpretability and controllability. For example, consider using disentangled representations to elucidate the diverse factors, such as style, brand, popularity, and price, that shape user preferences in dress selection. The inherent abstraction of latent factors from current disentangling methods makes it difficult to pinpoint which factor represents each specific dress attribute. In particular, one factor might be loosely related to a combination of brand and price, while another factor might represent a mixture of style and popularity. This ambiguity in the latent factors limits the interpretability of the recommendation system, making it difficult to understand why a particular item was recommended to a user. Moreover, this lack of clarity also affects the controllability of the recommendation system. Suppose a user wants to receive recommendations specifically for dresses without considering her preferences for price and popularity. It would be challenging to adjust the recommender system to focus on those specific preferences, as the latent factors are not clearly tied to these attributes.
@@
In summary, the contributions of this paper are threefold:
@@ 2.1. Multimodal Collaborative Filtering
Traditional Collaborative Filtering (CF) (Sarwar et al., 2001; Ebesu et al., 2018; Rendle et al., 2009) methods primarily rely on user-item interactions to learn representations of users and items. Consequently, the recommendation performance is negatively impacted when encountering users and items with limited interactions. To mitigate the challenges posed by data sparsity (McAuley and Leskovec, 2013; Kalantidis et al., 2013; Tan et al., 2016; Liu et al., 2023b), recent research has incorporated multimodal information, which has been extensively explored in other fields (Xu et al., 2015; Li et al., 2023b, 2024, a), into recommendation systems. The multimodal features (He and McAuley, 2016; Zhang et al., 2017a; Liu et al., 2019; Wei et al., 2019, 2020; Liu et al., 2023a; Du et al., 2023; Guo et al., 2024), such as reviews and images, provide valuable information for user preference and item characteristics, which can supplement historical user-item interactions and thus enhance recommendation performance.
Previous studies integrate multimodal information into the matrix factorization-based method in a straightforward manner. For instance, VBPR (He and McAuley, 2016) directly concatenates the visual features learned from item images with collaborative features as the joint item representation and feeds it into the MF module. With the success of deep learning techniques in modeling complex interaction behaviors (He et al., 2017) and the relation between various multimodal features (Nie et al., 2016), a lot of deep learning techniques have been introduced into multimodal recommendation (He et al., 2017; Tay et al., 2018). For example, MAML (Liu et al., 2019) first fuses the item’s multimodal features and user features, then feeds it into an attention module to capture users’ diverse preferences. VECF (Chen et al., 2019) constructs a visually explainable collaborative filtering model using a multimodal attention network, facilitating the integrated coupling of diverse feature modalities. More recently, graph convolutional networks (GCNs) (Hamilton et al., 2017) have demonstrated their power capability of representation learning for recommendation (Wang et al., 2019; He et al., 2020; Zhou et al., 2023; Wei et al., 2023). Based on the GCN structure, MMGCN (Wei et al., 2019) constructs a user-item bipartite graph to learn representations of each modality and then fuse the modality representation together as the final representation. GRCN (Wei et al., 2020) utilizes the rich multimodal content of items to refine the structure of the interaction graph in order to mitigate the effect of false-positive edges on recommendation performance.
@@ 2.2. Disentangled Representation Learning
Disentangled representation learning, which seeks to identify and separate underlying explanatory factors within data, has garnered significant interest, especially in the field of computer vision (Higgins et al., 2017; Liu et al., 2020; Ge et al., 2021). For instance, the beta-VAE method (Higgins et al., 2017) employs a constrained variational framework to learn disentangled representations of fundamental visual concepts. Meanwhile, IPGDN (Liu et al., 2020) automatically uncovers independent latent factors present in graph data.
Owing to the successful application of disentangled representation learning in various domains, numerous studies have concentrated on learning disentangled representations for users and items in recommendation systems in recent years (Ma et al., 2019; Tran and Lauw, 2022; Wang et al., 2020; Mu et al., 2022; Liu et al., 2023a; Wang et al., 2023). Following the previous work in computer vision, the initial attempts take Variational Auto-encoder (VAE) (Kingma and Welling, 2014) to learn disentangled representations. For example, MacridVAE (Ma et al., 2019) captured user preferences regarding the different concepts associated with user intentions separately. Beyond the disentangled representations derived from user-item interaction modeling, ADDVAE (Tran and Lauw, 2022) acquires an additional set of disentangled user representations from textual content and subsequently aligns both sets of representations. For studying the diversity of user intents on adopting the items, DGCF (Wang et al., 2020) employs a graph disentangled module to iteratively refine the intent-aware interaction graph and factorial representations for recommendations. In order to model the users’ various preferences on different factors of each modality, DRML (Liu et al., 2023a) estimates the user’s attention weight to underlying factors of different modalities with an attention module. As the learned disentangled representations lack clear meanings, KDR (Mu et al., 2022) uses knowledge graphs (KGs) to guide the learning process, ensuring that the resulting disentangled representations are associated with semantically meaningful information extracted from the KGs.
@@ 3.1.1. Problem Setting
Given a set of users $\mathcal{U}\in\{u\}$ and items $\mathcal{I}\in\{i\}$, we utilize three types of information to learn user and item representations: (1) A user-item interaction matrix $\bm{R}$, where each entry $r_{ui}\in\bm{R}$ represents the implicit feedback (e.g., clicks, likes or purchases) of user $u$ to item $i$. $r_{ui}=1$ indicates that user $u$ has interacted with item $i$ and $r_{ui}=0$ indicates that there is no interaction between user $u$ and item $i$ in the observed data; (2) Multimodal features, which mainly consist of two types of features associated with items, namely reviews and images; and (3) Attribute information, consisting of $K$ attributes and their associated attribute values of items, used as supervision for disentangled representation learning. We aim to learn robust representations of users and items, thereby predicting a user’s preference for items they have not yet interacted with.
@@
Similar to previous work (Wei et al., 2020; Liu et al., 2023a), each user $u$ and item $i$ are assigned with a unique ID and respectively represented by a vector $\bm{v}_{u}\in\mathbb{R}^{d}$ and $\bm{v}_{i}\in\mathbb{R}^{d}$, which are randomly initialized in our model. For the review and image information, we use the BERT (Devlin et al., 2019) and the ViT (Dosovitskiy et al., 2021) to extract the raw textual features $\bm{e}_{t}\in\mathbb{R}^{d_{0}}$ and visual features $\bm{e}_{v}\in\mathbb{R}^{d_{0}}$, respectively. To tailor the recommendation-oriented features, we adopt two non-linear transformations to cast $\bm{e}_{t}$ and $\bm{e}_{v}$ into the same feature space:
(1) $\bm{v}_{t}$ $=\sigma(\bm{W}_{t}\bm{e}_{t}+\bm{b}_{t}),$
$\bm{v}_{v}$ $=\sigma(\bm{W}_{v}\bm{e}_{v}+\bm{b}_{v}),$
where $\bm{W}_{t}$, $\bm{W}_{v}\in\mathbb{R}^{d\times d_{0}}$ and $\bm{b}_{t}$, $\bm{b}_{v}\in\mathbb{R}^{d}$ denote the weight matrices and bias vectors for textual and visual modality, respectively. $\sigma(\cdot)$ is the activation function.
Following the disentangled representation learning process in previous work (Wang et al., 2020; Liu et al., 2023a), we split the feature vector into $K$ chunks, each corresponding to a specific item attribute (e.g., price, brand). For simplicity, we equally split each modality’s representation into $K$ continuous chunks. Take item ID embedding as an example:
(2) $\bm{v}_{i}=(\bm{v}_{i}^{1},\bm{v}_{i}^{2},\cdots,\bm{v}_{i}^{K}),$
where $\bm{v}_{i}^{k}\in\mathbb{R}^{\frac{d}{K}}$ is the item ID embedding corresponding to the $k$-th attribute. Analogously, $\bm{v}_{t}=(\bm{v}_{t}^{1},\bm{v}_{t}^{2},\cdots,\bm{v}_{t}^{K})$, $\bm{v}_{v}=(\bm{v}_{v}^{1},\bm{v}_{v}^{2},\cdots,\bm{v}_{v}^{K})$ and $\bm{v}_{u}=(\bm{v}_{u}^{1},\bm{v}_{u}^{2},\cdots,\bm{v}_{u}^{K})$ are defined as the embeddings of textual feature, visual feature and user ID embedding, respectively.
@@ 3.1.3. Intuition of Attribute-driven Disentanglement
Our work not only aims to recommend accurate items to users but also advocates for attribute-driven disentanglement in multimodal recommender systems to enhance the interpretability and controllability of recommendations. Specifically, unlike the existing method that disentangles the latent factors in an unsupervised manner, we leverage the semantic labels of item attributes to learn attribute-specific sub-spaces for each attribute to obtain disentangled representations. In this paper, the vector of each modality is composed of $K$ chunks, with the assumption that each chunk is associated with an attribute (such as price or brand). The achievement of attribute-driven disentangling requires two necessary conditions. Firstly, each chunk should have a clear attribute reference, and chunks associated with different attributes should be distinguishable. For example, chunk $j$ is associated with the price, while chunk $k$ represents the vector of the brand. Secondly, each chunk of a specific attribute should accurately capture the specific attribute value. For example, chunk $j$ should reflect the price level (i.e., expensive or cheap) of the product.